OpenStudy (anonymous):

h(t)=-15t^2+105t+10,0

OpenStudy (anonymous):

can someone help me please

OpenStudy (cwrw238):

190 =-15t^2+105t+10 rearranging 15t^2 - 105t + 180 = 0 you need to solve this equation for t

OpenStudy (anonymous):

how do I do that

OpenStudy (cwrw238):

ok first take out the common factor 15 to simplify the problem: 15(t^2 -7t + 12) = 0 now we have to solve the trinomial in the parentheses we need 2 numbers whose product id +12 and whose sum is -7 these are -4 and -3 so it factors to ( t - 4)(t - 3) = 0 ginig t = 3 and t = 4 so your required times are 3 seconds and 4 seconds

OpenStudy (anonymous):

so the tshirt will reach the fan at 3 seconds

OpenStudy (cwrw238):

yea and again at 4 seconds

OpenStudy (cwrw238):

unless it gets sucked into the fan in 3 seconds!

OpenStudy (cwrw238):

if so its just as well its a tshirt